# Probing Einstein-Maxwell-Scalar Black hole via Thin Accretion Disks and Shadows with EHT Observations of M87* and Sgr A*

### v2

@@ III.1 Constraints on parameters with the EHT observations of M87* and Sgr A*
In this subsection, we find constraints to the coupling parameters $\alpha$ and $\beta$ using the observation data provided by the EHT for M87* and Sgr A*. We only focused on the non-rotating case since the rotation parameter of Sgr A* is small enough to have considerable deviation to the shadow radius 110 . Furthermore, it has also been concluded for M87* that it is difficult to distinguish between Kerr and dilaton BHs based on BH shadow images alone using general-relativistic magnetohydrodynamical simulations and radiative-transfer calculations to generate synthetic shadow images in comparison to the present observation from the EHT 111 . Also remarked in Ref.112 that the shadow size of M87* lies within the range of $3\sqrt{3}(1\pm 0.17)M$, whether their model is spherically symmetric or axisymmetric.
[…]
$d_{\mathrm{sh}}=\frac{D\theta}{M}$ (25)
Hence, the theoretical shadow diameter, however, can be obtained via
@@
$\left\{\begin{split}&d_{\mathrm{sh}}^{\mathrm{M}87^{*}}=(11\pm 1.5)M,\\ &d_{\mathrm{sh}}^{\mathrm{Sgr}.\mathrm{A}^{*}}=(9.5\pm 1.4)M.\\ \end{split}\right.$ (27)
M87*$/\alpha$ $2\sigma$ $2\sigma$ $1\sigma$ $1\sigma$ M87*/$\beta$ $2\sigma$ $2\sigma$ $1\sigma$ $1\sigma$
@@
Sgr A*$/\alpha$ $2\sigma$ $2\sigma$ $1\sigma$ $1\sigma$ Sgr A*/$\beta$ $2\sigma$ $2\sigma$ $1\sigma$ $1\sigma$
@@
The variation of the diameter of the shadow image with coupling parameter $\alpha$ and $\beta$ for M87* and for Sgr A* is shown in Fig.4 , showing uncertainties at $1\sigma$ and $2\sigma$ levels. The numerical values for the upper or lower bounds in $\beta$ and $\alpha$ is found in Table I and Table II. However, it turns out that the data for M87* gives a better constraint for the coupling parameter $\beta$. It would mean that there is a certain value for the coupling parameter that gives the observed value of M87* shadow. For the case of Sgr A*, we find that the parameter $\alpha$ remains unconstrained within the $1\sigma$ and $2\sigma$ confidence intervals, whereas the parameter $\beta$ is well constrained within the $1\sigma$ interval.
@@ III.2 Light bending: direct emission, lensing ring and photon ring
To gain a more comprehensive understanding of the appearance of EMS BH, we analyze its photon rings, and lensing rings in the presence of an optically bright accretion disk 177 ; 178 ; 179 . We begin by investigating the trajectory of light rays propagating in the vicinity of the EMS BH. To facilitate this analysis, we introduce the transformation $u=1/r$, which reformulates the orbit equation as follows
$(\frac{du}{d\phi})^{2}=G(u),$ (28)
where
@@
For a static and spherically symmetric spacetime, we consider the motion of a particle confined to the equatorial plane. From the Lagrangian, two conserved quantities can be obtained from Eq. (10).
The impact parameter is defined as $b\equiv\frac{p_{t}}{p_{\phi}}$. When $b>b_{c}$, a light ray originating from infinity approaches the black hole, reaches a closest approach, and then escapes back to infinity. For $b<b_{c}$, the light ray inevitably falls into the black hole. In the critical case $b=b_{c}$, the light ray asymptotically orbits the black hole at the photon sphere radius $r_{\mathrm{ph}}$.
For $b>b_{c}$, the turning point of the trajectory corresponds to the smallest positive real root of $G(u)=0$, denoted as $u_{m}$. According to Eq. (28), the total change in the azimuthal angle $\phi$ along a trajectory with a given impact parameter $b$ can be expressed as
$\phi=2\int_{0}^{u_{m}}\frac{du}{\sqrt{G(u)}},\quad b>b_{c}.$ (30)
For $b<b_{c}$, we only focus on the trajectory outside the horizon, so the total change of azimuthal angle $\phi$ is obtained by
$\phi=\int_{0}^{u_{0}}\frac{du}{\sqrt{G(u)}},\quad b<b_{c}$ (31)
where $u_{0}=1/r_{+}$.
Figure 5: The behavior of photon trajectories around the Schwarzschild, RN, and EMS black holes as a function of the impact parameter b. In the upper panel, we present the total number of orbits, defined as $n=\phi/2\pi$. The trajectories are categorized based on $n$, where direct emission $n<3/4$ is shown in black, lensed trajectories 3/$43/4<n<5/$ 4 in yellow, and photon ring trajectories $n>5/4$ in red. The lower panel displays selected photon trajectories in Euclidean polar coordinates $(r,\phi)$. The spacing in the impact parameter is set to $1/10$, $1/100$, and $1/1000$ for the direct, lensed, and photon ring trajectories, respectively. The BH is represented as a solid disk, while the dashed black circle in the ray-tracing diagram marks the photon orbit. For the three cases studied, we set $\alpha=0,\beta=0,q=0$ for the Schwarzschild BH (first column), $\alpha=0,\beta=0,q=0.5$ for the RN BH (second column), and $\alpha=0.5,\beta=0.8,q=0.5$ for the EMS BH (third column).
To analyze the observational signatures of radiation emitted from the vicinity of a BH accretion disk. The Ref.117 divide trajectories into direct, lensed and photon rings. One define the total number of orbits $n=\frac{\phi}{2\pi}$ which is obviously a function of impact parameter $b$. We denote the solution of
$n(b)=\frac{2m-1}{4},\quad m=1,2,3,\cdots$ (32)
by $b_{m}^{\pm}$. Note that $b_{m}^{-}<b_{c}$ and $b_{m}^{+}>b_{c}$. Then we can classify all trajectories as follows
@@
The physical picture of this classification is clear from the trajectory plots in Fig.5. Assuming light rays emit from north pole direction (far right of the trajectory plots), trajectories whose number of orbits 1/$41/4<n<3/$ 4 will intersect the equatorial plane only once. Trajectories whose number of orbits 3/$43/4<n<5/$ 4 will intersect the equatorial plane twice. Trajectories whose number of orbits $n>5/4$ will intersect the equatorial plane at least 3 times.
@@ III.3 Transfer functions and observed specific intensities
We analyze the emitted intensity from the EMS BH, assuming that the disk radiates isotropically in the rest frame of a static observer. According to Liouville’s theorem, the quantity $I_{\nu}^{\mathrm{em}}/\nu_{e}^{3}$ remains conserved along the trajectory of the light ray. Consequently, the observed intensity can be expressed as
@@
[…]
@@
$I_{\mathrm{em}}$ here is the total emitted specific intensity from the accretion disk. Therefore, the total intensity received by the observer will be
@@
where $r_{m}(b)$ is the $m$ intersection outside the horizon in the equatorial plane, which we call the transfer function. The transfer function establishes the connection between the radial coordinates and the photon’s impact parameter. It is important to note that our analysis does not account for light absorption or reflection by the accretion disk, nor does it consider the attenuation of light intensity due to environmental effects. Instead, we work within an idealized model.
We denote the solution of the orbit equation by $u(\phi,b)$ and focus on the first three transfer functions, which can be obtained by
$\left\{\begin{split}&r_{1}(b)=\frac{1}{u\left(\frac{\pi}{2},b\right)},&b\in\left(b_{1}^{-},\infty\right)\\ &r_{2}(b)=\frac{1}{u\left(\frac{3\pi}{2},b\right)},&b\in\left(b_{2}^{-},b_{2}^{+}\right)\\ &r_{3}(b)=\frac{1}{u\left(\frac{5\pi}{2},b\right)},&b\in\left(b_{3}^{-},b_{3}^{+}\right)\end{split}\right.$ (36)
As described in Ref.120 , the first transfer function corresponds to the “direct image” of the disk, which primarily reflects the redshift of the source profile. The second transfer function produces a highly demagnified image of the disk’s far side, known as the “lensing ring”. The third transfer function yields an extremely demagnified image of the near side of the disk, referred to as the “photon ring”. Images generated by subsequent transfer functions are so severely demagnified that they can be effectively disregarded.
@@
In Fig.6, the black dots $(m=1)$ represent the transfer function for direct emission and exhibit an almost constant slope, indicating a redshifted source profile. The yellow dots $(m=2)$ correspond to the lensing ring, where the slope remains small near $b_{c}$ but increases rapidly with $b$, leading to significant demagnification of the image from the far side of the accretion disk. The red dots $(m=3)$ represent the photon ring, where the slope approaches infinity, implying extreme demagnification of the disk’s near-side image. Consequently, the dominant contribution to the observed flux arises from direct emission, while higher-order images $(m>3)$ contribute negligibly and can be disregarded.
All three black holes exhibit these characteristics. The critical impact parameters are $b_{c}=5.19615$ for the Schwarzschild BH, $b_{c}=4.96791$ for the RN BH, and $b_{c}=4.76598$ for the EMS BH. This result indicates that the Schwarzschild BH has the largest critical impact parameter, corresponding to the largest shadow radius, while an increase in the model parameters leads to a reduction in the BH’s shadow size.
@@ III.4 Observational features of direct emission, photon and lensing rings
As discussed in the previous section, for an observer at infinity, the specific intensity depends solely on the radial coordinate $r$. In our analysis, we consider three toy models for the emissivity profile $I^{\mathrm{em}}$.
With the transfer function determined, we now consider a specific emission profile. As the first model, we examine a scenario in which the emission is sharply concentrated at the innermost stable circular orbit (ISCO) given by
@@
Second model, we consider the emission is sharply peaked at the photon sphere, and it ends abruptly at $r=r_{\text{ph}}$ while quickly decaying to zero, such as
@@
Final model, we consider an emission decaying gradually from the horizon to the ISCO, such as
@@
Each of these three models exhibits distinct characteristics. In particular, the second model features a rapid decay rate, whereas the third model exhibits a much slower decay. In the third model, the emission originates directly from the event horizon, while in the second model, it begins at the photon sphere, and in the first model, it is confined to the ISCO. Although these models represent highly idealized scenarios, they provide valuable qualitative insights into the behavior of photons around the BH.
@@
In the Fig.7, 8, and 9, we have shown the observed appearance of the thin accretion disk with the observed intensity and the impact factor corresponding to the three models for different coupling constant parameters respectively. In all the figures, we see that for the first model (top row), the emission intensity (top-first) has a peak near the critical impact parameter $r\sim b_{c}$ and then it decreases as radial distance increases and becomes zero. In this case, the photon sphere lies in the interior region of the emission part of the disk. As can be readily seen from the middle panel (top-second), the direct image of the disk looks very similar to the emission profile, although its abrupt end occurs at different place due to gravitational lensing. However, we noticed that the peak of the photon and lensing rings are not only smaller than the direct emission but also have a narrow observational area. Hence, we observed that the observed intensity has a huge contribution coming from the direct emission, a small contribution from the lensing rings, and very little contribution from the photon rings. As is evident from the right panel (2D shadow image, top-third) where the lensing ring emission appears as a tiny ring inside the direct image and one can barely see the photon ring emission in the right panel if one zooms in.
For the second model (middle row), the intensity of the emission has a peak at photon sphere $r\sim r_{\mathrm{ph}}$ and then decreases as radial distance increases (middle-first). In this case, the observed intensity profile (middle-second) peaks due to the direct emission, and then it shows an attenuation with increasing $r$. We observed that the photon and lensing rings, which improves the total intensity of this particular area, and therefore, we get a new peak due to the photon ring, lensing ring, and direct emission. However, the photon and lensing rings are highly demagnetized and have a very narrow area in the observed intensity. Therefore, we still have a dominant contribution coming from the direct emission in the observed intensity which can be seen in the $2\mathrm{D}$ shadow image (middle-third).
For the third model (bottom row), the peak in the intensity starts from the horizon ($r_{+}$) and then decreases with increasing $r$ (bottom-first). In this case, redshift effects noticeably decrease the observed flux. However, the most important difference from the top two row for our considerations is that the lensing ring and photon ring emission are now superimposed on the direct emission. However, as discussed in the last section, the photon and lensing rings are highly demagnetized and have a very narrow area in the observed intensity. Therefore, we still have a dominant contribution coming from the direct emission in the observed intensity which can be seen in the 2D shadow image (bottom-third).

### v3

@@ III.1 Constraints on parameters with the EHT observations of M87* and Sgr A*
We will obtain constraints on the coupling parameters $\alpha$ and $\beta$ using EHT data for M87 and Sgr A, focusing on the non-rotating case. This simplification is justified because
  • M87: General-relativistic magnetohydrodynamic simulations indicate that current EHT observations cannot reliably distinguish between Kerr and dilaton BHs based on shadow morphology alone 111 . The observed shadow size is consistent with $3\sqrt{3}(1\pm 0.17)M$, regardless of spherical or axisymmetric models 112 .
  • Sgr A: The rotation parameter is sufficiently small to negligibly affect the shadow radius 110 .
Based on observational data, the shadow of M87* exhibits an angular diameter of $42\pm 3\ \mathrm{\rm\mu as}$, corresponding to a distance of $16.8$ Mpc from Earth and a central BH mass of $(6.5\pm 0.90)\times 10^{9}\ M_{\odot}$ 113 . In parallel, the EHT’s recent findings for Sgr A* report a shadow size of $48.7\pm 7\ \mathrm{\mu as}$, with the source located at a distance of $8277\pm 33$ pc and hosting a supermassive BH of mass $(4.3\pm 0.013)\times 10^{6}\ M_{\odot}$ 114 ; 115 .
The shadow diameter in mass units can be calculated from the derived BH parameters using the following expression 116
$d_{\mathrm{sh}}=\frac{D\theta}{M}.$ (26)
Hence, the theoretical shadow diameter can be obtained via
@@
[…]
Our constraints on the model parameters are shown in Fig.4. The corresponding numerical constraints are further presented in Tab.I (for M87*) and Tab.II(for Sgr A*).
[…]
@@
[…]
@@
The variation of the diameter of the shadow image with coupling parameter $\alpha$ and $\beta$ for M87* and for Sgr A* is shown in Fig.4, showing uncertainties at $1\sigma$ and $2\sigma$ levels. The numerical values for the upper or lower bounds in $\beta$ and $\alpha$ is found in Tab.I and Tab.II. However, it turns out that the data for M87* gives a better constraint for the coupling parameter $\beta$. It would mean that there is a certain value for the coupling parameter that gives the observed value of M87* shadow. For the case of Sgr A*, we find that the parameter $\alpha$ remains unconstrained within the $1\sigma$ and $2\sigma$ confidence intervals, whereas the parameter $\beta$ is well constrained within the $1\sigma$ interval.
@@ III.2 Light bending: direct emission, lensing ring and photon ring
To systematically investigate the observational characteristics of EMS BH, we examine their photon ring and lensing ring structures under the assumption of an optically thick accretion disk 177 ; 178 ; 179 . Our analysis begins with studying null geodesics in the EMS BH.
For computational convenience, we employ the coordinate transformation $u\equiv 1/r$, which yields the compact orbital equation
$\left(\frac{du}{d\phi}\right)^{2}=G(u),$ (29)
where the function $G(u)$ takes the explicit form
@@
Similarly, we restrict our analysis to equatorial motion. The system’s Lagrangian formalism provides two conserved quantities, as derived in Eq.(13). The critical impact parameter $b_{c}$ separates distinct null geodesic behaviors
  • Case 1 ($b>b_{c}$): Photons from asymptotic infinity are gravitationally scattered, reaching a minimum coordinate distance $r_{\text{min}}$ before returning to infinity.
  • Case 2 ($b<b_{c}$): Photons become captured by the BH, demonstrating the existence of a photon capture cross-section.
  • Case 3 ($b=b_{c}$): Photons spiral asymptotically toward the photon sphere at $r_{\text{ph}}$, defining the boundary of BH shadow.
For $b>b_{c}$, the turning point of the trajectory corresponds to the smallest positive real root of $G(u)=0$, denoted as $u_{\rm m}=1/r_{\rm min}$ with $r_{\rm min}$ being the light ray’s radial minimal distance from its trajectory to the BH. Hence, the total change in the azimuthal angle $\phi$ outside the horizon of the trajectory with a given impact parameter $b$ can be expressed as 178 ; 179
$\phi=2\int_{0}^{u_{\rm m}}\frac{du}{\sqrt{G(u)}},\quad b>b_{c}.$ (31)
For $b<b_{c}$, the total change of angle $\phi$ is given by
$\phi=\int_{0}^{u_{0}}\frac{du}{\sqrt{G(u)}},\quad b<b_{c},$ (32)
where $u_{0}\equiv 1/r_{+}$.
Figure 5: The behavior of photon trajectories around the Schwarzschild, RN, and EMS BHs as a function of the impact parameter $b$. In the upper panel, we present the total number of orbits, defined as $n=\phi/2\pi$. The trajectories are categorized based on $n$, where direct emission $n<3/4$ is shown in black, lensed trajectories $3/4<n<5/4$ in yellow, and photon ring trajectories $n>5/4$ in red. The lower panel displays selected photon trajectories in Euclidean polar coordinates $(r,\phi)$. The spacing in the impact parameter is set to $1/10$, $1/100$, and $1/1000$ for the direct, lensed, and photon ring trajectories, respectively. The BH is represented as a solid disk, while the dashed black circle in the ray-tracing diagram marks the photon orbit. For the three cases studied, we set $\alpha=0,\beta=0,q=0$ for the Schwarzschild BH (first column), $\alpha=0,\beta=0,q=0.5$ for the RN BH (second column), and $\alpha=0.5,\beta=0.8,q=0.5$ for the EMS BH (third column).
Following the trajectory classification established in Refs. 120 ; 178 where photon orbits are categorized into direct, lensed, and photon ring components, we quantify the orbital winding through the turn number $n(b)\equiv\phi/2\pi$. This quantity exhibits discrete quantization as demonstrated in Refs. 178 ; 179 , with the functional dependence on impact parameter $b$ given by
$n(b)=\frac{2m-1}{4},\quad m\in\mathbb{Z}^{+}.$ (33)
The critical impact parameters $b_{m}^{\pm}$ demarcate these orbital families, where $b_{m}^{-}$ (satisfying $b_{m}^{-}<b_{c}$) and $b_{m}^{+}$ (satisfying $b_{m}^{+}>b_{c}$) correspond respectively to the minimal and maximal solutions for each $m$-th order orbit, with $b_{c}$ denoting the photon sphere critical parameter. Then the rays can be classified as follows
@@
The behaviors of photons in different BHs are plotted in Fig.5. The up panel of Fig.5 depicts the total number of orbits. And the down panel depicts the trajectories of light rays surrounding the BHs, such as ”direct” will intersect the equatorial plane only once.
@@ III.3 Transfer functions and observed specific intensities
In this subsection, we will derive the specific intensity $I_{\nu}^{\mathrm{em}}$ radiated by the accretion disk, assuming local isotropy of radiation in the rest frame of static observers. The observed intensity can be expressed as 177 ; 178 ; 179
@@
where $g=\sqrt{f(r)}$ and $I_{\nu^{\prime}}^{obs}$ are the observed intensity at the frequency $\nu^{\prime}$. By integrating over all frequencies, we obtain the total emitted intensity as $I^{\mathrm{em}}=\int I_{\nu}^{\mathrm{em}}d\nu$. Thus, the corresponding observed frequency can be written as
@@
$I^{\mathrm{em}}$ denotes the integrated specific intensity radiating from the entire accretion disk. Hence, the total received intensity is
@@
where $r_{m}(b)$ is transfer function (It represents the $m$ intersection points with the equatorial plane). The transfer function fundamentally connects radial coordinates with photon impact parameters, we emphasize that our analysis operates under two key idealizations: (1) no light absorption or reflection by the disk, and (2) no environmental intensity attenuation.
Therefore, the first three transfer functions, which can be obtained by 177 ; 178 ; 179
$\left\{\begin{split}&r_{1}(b)=\frac{1}{u\left(\frac{\pi}{2},b\right)},&b\in\left(b_{1}^{-},\infty\right),\\ &r_{2}(b)=\frac{1}{u\left(\frac{3\pi}{2},b\right)},&b\in\left(b_{2}^{-},b_{2}^{+}\right),\\ &r_{3}(b)=\frac{1}{u\left(\frac{5\pi}{2},b\right)},&b\in\left(b_{3}^{-},b_{3}^{+}\right).\end{split}\right$. (37)
As established in Ref.120 , three distinct observational features emerge:(1) The first transfer function corresponds to the direct disk image;(2) The second forms a significantly demagnified lensing ring from the disk’s far side;(3) The third creates the extremely faint photon ring from the near side. The corresponding graphical representation is presented in the figure below.
@@
As shown in Fig.6, the black points $(m=1)$ correspond to the direct emission transfer function and exhibit an approximately constant slope, reflecting the redshifted emission profile from the disk. The yellow points $(m=2)$, associated with the lensing ring, display a rapidly increasing slope with growing impact parameter $b$, indicating that the far-side image is significantly demagnified. The red points $(m=3)$ denote the photon ring, where the slope becomes nearly infinite, resulting in extreme demagnification of the near-side image. Thus, the observed flux is primarily governed by the direct emission, with higher-order contributions $(m>3)$ being negligible.
All three BHs exhibit these characteristics. The critical impact parameters are $b_{c}=5.19615$ for the Schwarzschild BH, $b_{c}=4.96791$ for the RN BH, and $b_{c}=4.76598$ for the EMS BH. This result indicates that the Schwarzschild BH has the largest critical impact parameter, corresponding to the largest shadow radius, while an increase in the model parameters leads to a reduction in the BH’s shadow size.
Additionally, we not only learn the aforementioned characteristics but can also quantify both the morphology (width) and intensity changes of the three ring types. First, we discuss how scalar couplings affect the photon ring. As shown in Table III, we have following calculations. In Schwarzschild BH ($q=0,\alpha=0,\beta=0$), we have $\Delta b_{\text{ph}}=b_{3}^{+}-b_{3}^{-}=5.22794-5.18781=0.04013$. In EMS BH ($\alpha=0.5,\beta=0.8,q=0.5$), we have $\Delta b_{\text{ph}}=b_{3}^{+}-b_{3}^{-}=4.80685-4.75392=0.05293$. The EMS BH shows a net increase of $32\%$ in width. A possible explanation is: the scalar field flattens the effective potential near $r_{\text{ph}}$ (due to $C(r)=r^{2}(1-b_{2}/r)^{\frac{2\alpha^{2}}{1+\alpha^{2}}}$), allowing more near-critical orbits.
For lensing ring adjustment, we now properly treat the two disjoint intervals ($b\in(b_{2}^{-},b_{3}^{-})$ and $b\in(b_{3}^{+},b_{2}^{+})$). In Schwarzschild BH ($q=0,\alpha=0,\beta=0$), we have $\Delta b_{\text{lens1}}=b_{3}^{-}-b_{2}^{-}=5.18781-5.01514=0.17267$, $\Delta b_{\text{lens2}}=b_{2}^{+}-b_{3}^{+}=6.16757-5.22794=0.93963$. Total $\Delta b_{\text{lens}}=1.1123$. In EMS BH ($\alpha=0.5,\beta=0.8,q=0.5$), we have $\Delta b_{\text{lens1}}=b_{3}^{-}-b_{2}^{-}=4.75392-4.55452=0.1994$, $\Delta b_{\text{lens2}}=b_{2}^{+}-b_{3}^{+}=5.81006-4.80685=1.00321$. Total $\Delta b_{\text{lens}}=1.20261$. Compared to a Schwarzschild BH, the EMS BH shows a net increase of $8.1\%$. While the total lensing ring width increases modestly, the inner sub-interval ($\Delta b_{\text{lens1}}$) shows a $15.5\%$ increase due to stronger light bending near the photon sphere.
Finally we consider the direct emission, increasing $\alpha$ leads to a larger $b_{c}$, whereas increasing $\beta$ decreases $b_{c}$. From Table III, the boundaries of the direct emission region exhibit a systematic shift along the impact-parameter axis.
@@ III.4 Observational features of direct emission, photon and lensing rings
As discussed in the previous section, for an observer at infinity, the specific intensity depends solely on the radial coordinate $r$. In our analysis, we consider the following three toy models for the emissivity profile $I^{\mathrm{em}}$177 ; 178 ; 179 ,
@@
where $r_{\text{isco}},r_{\text{ph }}$ and $r_{\mathrm{+}}$ is the innermost stable orbit, photon sphere and event horizon respectively. These three models each exhibit distinct characteristics. The following figures present the results for three distinct types of BHs.
@@
Figures 7, 8, and 9 display the observed emission from thin accretion disks, intensity distributions, and impact parameter mappings for three BH models under different coupling parameters. In each case, the top row corresponds to the first model. The intensity profile (top-left panel) peaks near the critical impact parameter $b_{c}$ and gradually decreases outward. The direct image (top-middle panel) closely follows the intrinsic emission but is affected by gravitational lensing, leading to a sharp boundary and image distortion. Notably, the photon and lensing rings contribute marginally to the overall intensity due to their limited angular extent and lower peak brightness. The shadow image (top-right panel) reveals the lensing ring as a narrow, bright ring just inside the direct image, while the photon ring appears extremely faint and confined to a narrow region, visible only upon magnification. These results highlight the dominant role of direct emission in observed disk images.
In the second model (middle row), the emission intensity peaks near the photon sphere ($r\sim r_{\mathrm{ph}}$) and gradually decreases outward (middle-left panel). The observed intensity profile (middle-center) similarly shows a peak dominated by direct emission from this region, followed by a steady decline. Although the photon and lensing rings contribute modestly, their overlap with direct emission slightly enhances the total signal. Nonetheless, due to strong gravitational redshift and depolarization, their brightness remains faint and confined to a narrow angle. As shown in the shadow image (middle-right panel), direct emission remains the dominant observable feature.
In the third model (bottom row), the emission intensity peaks near the event horizon ($r_{+}$) and gradually declines with increasing radius (bottom-left panel). Due to strong gravitational redshift, much of this radiation is suppressed before reaching a distant observer. Unlike the previous models, the photon and lensing ring emissions overlay the direct radiation; however, their contributions remain minimal because of severe demagnification and narrow angular extent. Consequently, the observed intensity is still primarily determined by direct emission, as clearly seen in the shadow image (bottom-right panel).
@@
III.5 Weak gravitational lensing and magnification of lensed image
Let us now examine the brightness of the image in the weak field limit by using the light’s deflection angle around EMS BH. We will find the deflection angle by using the Gauss-Bonnet theorem (GBT), originally stated as 667 ; 668
$\iint_{D}KdS+\sum_{a=1}^{N}\int_{\partial D_{a}}\kappa_{\mathrm{g}}d\ell+\sum_{a=1}^{N}\theta_{a}=2\pi\chi(D).$ (41)
Consider a freely orientable $2D$ curved surface $S$ with an infinitesimal area element $dS$, which defines a domain $D$ with a Gaussian curvature $K$. The boundary of $D$ is given by $N$ components, $\partial D_{\mathrm{a}}$ (where $\mathrm{a}=1,2,\ldots,N$), along which the geodesic curvature $\kappa_{\mathrm{g}}$ is integrated over the path element $d\ell$ following a positive convention. The jump angles are denoted by $\theta_{\mathrm{a}}$. The Euler characteristic of the domain, $\chi(D)$, is equal to 1 in this case since $D$ is a non-singular region.
It was shown by Ref.666 that in spherically symmetric spacetime admitting asymptotic flatness, the deflection angle can be written as
$\hat{\alpha}=-\iint_{D}K\,dS.$ (42)
Substituting the EMS solution into the above expression and performing the weak-field approximation yields
$\hat{\alpha}\approx\frac{4M}{b}-\frac{3\pi Q^{2}\left(1-\alpha^{2}+\beta\right)}{4b^{2}}.$ (43)
In the context of gravitational lensing, let us consider the following lens equation, expressed in terms of the light angles $\hat{\alpha}$, $\theta$, and $\eta$ Morozova13 ; Bozza2008lens ; Babar21a ; Al_Badawi_2024
$\theta D_{\mathrm{s}}=\eta D_{\mathrm{s}}+\hat{\alpha}D_{\mathrm{ds}}\,$, (44)
where $D_{\mathrm{s}}$ is the distance between the source and observer, $D_{\mathrm{ds}}$ is the distance between the source and lens, and the angular positions of the image and the source are denoted by $\theta$ and $\eta$, respectively. From Eq. (44), the source’s angular position can be expressed as
$\eta=\theta-\frac{D_{\mathrm{ds}}}{D_{\mathrm{s}}}\frac{\hat{\alpha}(\theta)}{D_{\mathrm{d}}}\,$, (45)
where $D_{\mathrm{d}}$ is the distance between the lens and observer. We note the use of the reduced deflection angle $\hat{\alpha}(\theta)=\xi(\theta)/b$, where $b\equiv D_{d}\theta$ is the impact parameter Bozza2008lens . The angular part, known as the Einstein radius $\theta_{E}$, is derived from the spacetime geometry between the source images and corresponds to the radius of the Einstein ring, provided the image forms a ring-like shape. It is given by Morozova13
$\theta_{E}=\sqrt{2R_{s}\frac{D_{ds}}{D_{d}D_{s}}}\,$, (46)
with $R_{s}=D_{d}\theta_{E}$ as the radius of the Einstein ring.
Next, we investigate the magnification of brightness, defined as the ratio of the total brightness of the lensed image ($I_{\mathrm{tot}}$) to the unlensed source brightness ($I_{*}$). The total magnification is given by Babar21a ; Atamurotov2022
$\mu_{\mathrm{tot}}=\frac{I_{\mathrm{tot}}}{I_{*}}=\sum_{k}\bigg|\frac{\theta_{k}}{\beta}\frac{d\theta_{k}}{d\beta}\bigg|,\quad k=1,2,\dots,j\,.$ (47)
For the specific case of two images, the total magnification simplifies to
$\mu_{\mathrm{tot}}=\frac{x^{2}+2}{x\sqrt{x^{2}+4}}\,$, (48)
where $x=\eta/\theta_{E}$ is a dimensionless parameter.
We conducted a numerical analysis of the total magnification for different values of the parameters $\alpha$, $\beta$, and $q$ for EMS BH (Fig  10). Our findings show that for a fixed value of $q$, the total magnification increases with increasing $\alpha$ but decreases with increasing $\beta$. Furthermore, the total magnification decreases as the BH charge increases. We also observed a downward shift in the magnification curves for increasing values of $\beta$ when $\alpha=0$. Conversely, for $\beta=0$, we obtain the opposite trend compared to the $\alpha=0$ case.
Figure 10: Total magnification $\mu_{tot}$ plotted as a function of the different parameters $\alpha$, $\beta$ and $q$.